OpenStudy (anonymous):

Find A and B given that the function y=Ax^(-1/2)+Bx^(1/2) has a minimum value of 6 at x = 9

myininaya (myininaya):

So you found y' correct?

OpenStudy (anonymous):

No I found \[B = (A/-9)+2\]

OpenStudy (anonymous):

Use \[ y(6)=9 \]and \[ y'(6)=0 \]Those two equations will allow you to solve for the variables.

OpenStudy (anonymous):

But after that I tried to plug in and then i didnt work out

OpenStudy (anonymous):

:O I didnt think about the derivative

OpenStudy (anonymous):

But wio didnt you get those backwards because x=9 and has a value of 6

OpenStudy (anonymous):

So its f(9)=6 and f'(9)=0

myininaya (myininaya):

You are right matt.

OpenStudy (anonymous):

Lol cool

OpenStudy (anonymous):

Ok so I solved for B and I pluged it all in and got this \[y=\frac{ A }{ \sqrt(x) }+(\frac{ -A }{ 9 }+2)\sqrt(x)\]

myininaya (myininaya):

How did you get B was equal to that?

OpenStudy (anonymous):

6=A(9)^(-1/2)+B(9)^(1/2) Since the sqrt(9)=3 It turns into this 6=(A/3)+B(3) Then I subtracted B(3) to the left side 6-B(3)=(A/3) Then I subtracted 6 on the right side -B(3)=(A/3)-6 Then I divided -3 on both sides to get B=(-A/9)+2

myininaya (myininaya):

Oh I left off the 3 in front of the B. Looks good.

myininaya (myininaya):

So we have one linear equation from the y(9)=6. now let's look at y'(9)=0.

myininaya (myininaya):

I would just leave in terms of A and B until the end. But you can do this like you are.

OpenStudy (anonymous):

Just start off by taking the derivative?

myininaya (myininaya):

yep.

OpenStudy (anonymous):

Oh well then lol

myininaya (myininaya):

A and B are just constants. \[(x^\frac{1}{2})'=? \text{ and } (x^\frac{-1}{2})'=?\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For some reason im at this point \[1=\frac{ A }{ 2x^3}-\frac{ A }{ 18 }\] But if I try to

OpenStudy (anonymous):

I dont know what im doing

myininaya (myininaya):

oh so did you keep it in terms of A and B.

myininaya (myininaya):

It doesn't matter. I just think it looks less confusing if we keep in terms of A and B but we can do in terms of A if you so wish.

OpenStudy (anonymous):

No its from where I left off but I guess ill try to start off with in terms of A or B because its confusing

myininaya (myininaya):

Is this what you got for y'? \[y'=\frac{-1}{2}A \frac{1}{x^\frac{3}{2}}+\frac{1}{2}(\frac{-A}{9}+2)\frac{1}{x^\frac{1}{2}}\]

myininaya (myininaya):

Did you do y'(9)=0?

myininaya (myininaya):

set x equal to 9 and y'=0?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ok now I got A=9

OpenStudy (anonymous):

So then B=1

OpenStudy (anonymous):

Correct?

myininaya (myininaya):

Yes. that looks good. :)

OpenStudy (anonymous):

Thanks a bunch myininaya your always a huge help! :D